OpenStudy (anonymous):

What does In mean? d/dxIn(12x+4) = ?

OpenStudy (anonymous):

I always get confused at what " In " means in an equation. Integral? Indefinite? :0

OpenStudy (anonymous):

it means "natural log"...you can just press the ln button on a calculator

OpenStudy (anonymous):

Lol my calculator ran out of battery so I have resulted to using wolfram but it doesn't seem to understand when I input "In" in the equation.

OpenStudy (accessdenied):

Are you sure it is an uppercase i or have you considered a lowercase L? (They look similar)

OpenStudy (accessdenied):

As in Ln, or Log / log base e.

OpenStudy (anonymous):

its a lowercase L

OpenStudy (anonymous):

Oh, I have never thought of it as Ln. I think thats why it wouldnt compute in wolfram. So it's "Ln" then?

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (accessdenied):

i wrote capital L to distinguish it from how it looks, but yes. That is called natural logarithm. Of course, by the fun of life lowercase L and uppercase i get to be strikingly similar to confuse. (and don't get me started on that vertical bar thing too, or 1)

OpenStudy (ipwnbunnies):

Absolute value irks you out?

OpenStudy (accessdenied):

mainly when I see it written out on paper, some people are just clumsy about it so you sometimes just can never tell! :p

OpenStudy (anonymous):

ln is just the natural logarithm 

OpenStudy (anonymous):

You don't need a calculator for this lol. d/dx of any ln expression is (expression)' * 1/(expression). So in your case it's (12x+4)' * 1/(12x+4) = 12/(12x+4).

OpenStudy (anonymous):

So would it be safe to put it as d/dx (log(12x+4) = 3/(3x+1) ?

OpenStudy (ipwnbunnies):

No, log(x) is different. It's just that the derivative of ln(x) plays out nice.

OpenStudy (anonymous):